# Are LLMs effective psychological assessors? Leveraging adaptive RAG for interpretable mental health screening through psychometric practice

### v1

@@ 4.3 Proposed Framework
The proposed workflow involves a two-step adaptive RAG pipeline. First, we retrieve the most relevant documents for each user with respect to each questionnaire item, and then we instruct the LLM to predict the corresponding score.
Figure 3: Results associated with the use of different combinations of LLMs, prompting strategies, and retrieval models across various metrics for the eRisk 2019 collection. The best combinations for each metric are highlighted with $\star$. The darker the color, the better the score.
In this work we explore various models for zero-shot retrieval and LLMs (both open- and closed-source) alongside different prompting techniques (Direct and CoT) to examine how the results vary across different combinations. All LLMs are used with temperature set to 0 to force deterministic outputs. For predictions, the closed-source models used include gpt-4o-mini and Claude-3.5-Sonnet, while the open-source models are Phi-3-mini and Phi-3.5-mini. For prompting, we use two techniques illustrated in Figure 2. The first technique generates the item scores directly based on relevant Reddit posts (Direct), while the second approach guides the LLM to reflect on intermediate steps (CoT), encouraging the LLM to go through reasoning steps before predicting the final score.
@@ 5 Results
eRisk 2020 BioInfo Oliveira (2020) 30.00% 76.01% 38.30% 69.21%
ILab Martínez-Castaño et al. (2020) 27.14% 81.70% 37.07% 69.41%
@@ 5.1 Predicting Psychological Questionnaire Scores
For evaluating the effectiveness of our approach in predicting responses to the BDI-II questionnaire, we use the official eRisk benchmark metrics Losada et al. (2019) which assess performance at two distinct levels (for all the considered metrics, the higher the value, the better):
At questionnaire level, we examine Depression Category Hit Rate (DCHR) which measures the accuracy in estimating depression severity levels (minimal: 0-9, mild: 10-18, moderate: 19-29, and severe: 30-63), and the Average Difference Between Overall Depression Levels (ADODL) which evaluates the overall BDI-II score estimations.
@@
The idea is to explore different combinations of LLMs, dense retrievers, and prompting strategies for the eRisk 2019 dataset, and using the best combinations for the eRisk 2020 collection. To answer (RQ.2), Figure 3 shows the results as heatmaps for each combination across the 4 metrics considered for this task for the eRisk 2019 dataset. On average, we observe that using a direct prompting technique leads to better results compared to a Chain of Thought (CoT) approach. Overall, we notice that closed-source LLMs often outperform open-source ones. Specifically, the worst results are obtained when using open-source models combined with CoT prompting techniques.
Table 4 shows the results of the best combinations for each metric for the corresponding eRisk edition, along with benchmark models. For comparison, we consider the best prior works for each metric for the eRisk 2019 and 2020 collections. We refer the reader to the corresponding overview for more in-depth details Losada et al. (2019, 2020). In Section C of the Appendix we further discuss all the models used as benchmarks. After filtering the best performing combinations from the Figure 3, these are also used for the 2020 edition. We observe that, even among the benchmark models used for comparison, there is no single model that outperforms others across all metrics. Regarding the 2019 edition, the proposed approaches outperform the benchmarks across all considered metrics, except for AHR, where only one combination manages to outperform that edition’s best model. For the 2020 edition, the comparison is more challenging because all the considered comparisons use the 2019 edition data for training. In particular, the recent work by Pérez et al. (2023), in a specific configuration, achieves SOTA results for questionnaire metrics, despite having low scores at the item level (AHR and ACR), where our combinations outperform the benchmarks. For depression category accuracy (DCHR), we obtain the best result with a voting-regressor ensemble of the top 3 models based on DCHR from 2019. This ensemble represents the rounded average of each prediction from the three models.
Interestingly, although the distribution of different categories changes between the two editions, the results, in terms of metrics, seem to show consistency between editions. No substantial performance drops are observed, except for DCHR. This negative result may be due to how overall scores are categorized into the 4 severity categories of the BDI-II questionnaire. Within the context of the challenge, the authors of the eRisk workshop use the BDI parameterization, the version preceding BDI-II. In BDI-II, new ranges are introduced that change from those of the previous test, especially regarding the minimum level of depression. Specifically, minimal or absent depression is identified as 0-13, Mild as 14-19, Moderate as 20-28, and severe as 29-63. Table 5 shows how DCHR changes when using the correct reparameterization, obtaining excellent and completely counterintuitive results, especially for the two models using Claude and the CoT strategy, compared to those obtained with the previous questionnaire parameterization in the 2020 dataset.
Collection Model P R F1
eRisk 2019 UNSL 0.71 0.41 0.52
iLab 0.68 0.66 0.67
OURS 0.65 0.61 0.62
eRisk 2020 iLab 0.83 0.69 0.75
OURS 0.90 0.52 0.67
@@
Table 6: Performance of our approach compared to the best approach in each of the two eRisk editions for the early detection of self-harm task. Precision, Recall, and F1 metrics are reported.
Collection Model $\mathbf{\tau}$ P R F1
@@
5.2 Identifying Signs of Self-Harm through questionnaire
We approach the identification of self-harm behaviors using the Self-Harm Inventory (SHI), a 22-item, yes/no self-report questionnaire, that screens for lifetime history of self-harm behaviors. A score of 5 or more “yes” responses on the SHI indicates potential mild forms of Deliberate Self-Harm (DSH) Latimer et al. (2009). To answer (RQ.3), the methodology follows the same approach used for BDI-II, with the key distinction that SHI is structured as a binary questionnaire, unlike BDI-II Likert scale. The process involves retrieving the most relevant documents using SHI questions as queries, followed by LLM-generated responses. For this specific task, we chose the combination of gpt-4o-mini with Direct Prompt and MiniLM-L12-v3 as the retrieval model, as this configuration demonstrated the best trade-off between performance, computational costs, and processing speed in previous tests. The dataset provides each user complete post history, associated with a binary label indicating the presence or absence of self-harm behaviors. While the original task aims to identify self-harm cases as early as possible using the minimum number of posts, our approach considers the entire history to maximize prediction accuracy. For the 2019 edition, we compared our approach with two reference models: UNSL Burdisso et al. (2019), which analyzes only a subset of posts, and iLab Martínez-Castaño et al. (2020), which uses BERT fine-tuned in a fully supervised context. iLab approach was optimized to maximize F1-score and trained on a custom dataset built from self-harm subreddit posts. To ensure a fair comparison, we considered the version of iLab that has access to the complete post history for both the 2019 and 2020 editions. Table 7 shows that while the fine-tuned BERT-based model, benefiting from an extensive training corpus, generally outperforms our unsupervised approach, the performance gap is remarkably narrow, particularly for the 2019 edition. Notably, our model achieves superior precision in the 2020 edition. These results are particularly significant considering that our approach is completely unsupervised and requires no training data, yet achieves competitive performance compared to fully supervised models that leverage extensive domain-specific training. This suggests that our unsupervised approach based on adaptive RAG can offer a viable alternative in scenarios where labeled training data is scarce or unavailable.
5.3 The Importance of Questionnaire for Screening Procedure
@@
For this task, we used data from the 2017 eRisk edition to address (RQ4.). As in the previous subsection, we have access to users post histories along with corresponding labels indicating whether each user exhibited clinical signs of depression.
@@
We tested the same adaptive RAG approach that proved effective for previous tasks, using the same combination of methods but guided by different types of standardized depression screening questionnaires. Specifically, we compared results obtained from the Beck Depression Inventory-II (BDI-II), Patient Health Questionnaire-9 (PHQ-9), Depression Anxiety Stress Scales-42 (DASS-42) (specifically focusing on its depression subscale), and simply instructing the LLM to determine if the patient was suffering from depression.
@@
Each of these three questionnaires has established optimal cut-off scores for identifying clinically significant depression, shown as $\tau$ in Table 7, which also presents the results from the different approaches. The worst results were obtained through direct instruction to gpt-4o-mini and through the use of the DASS-42 depression subscale. The best results in terms of precision were achieved using PHQ-9 (0.69), while BDI-II showed the highest recall (0.76). The highest F1 score (0.60) was achieved using an ensemble classifier combining all three questionnaires. To answer (RQ4.), our findings suggest that structured psychological assessment tools, can enhance the effectiveness of LLM-based mental health assessments compared to direct questioning approaches.
@@ 6 Conclusions
Our work demonstrates that LLMs, when guided by standardized psychological questionnaires through our novel adaptive RAG approach, can effectively support mental health screening tasks in a completely unsupervised setting. The results show the goodness of this approach in automatically completing questionnaires (RQ1.), proving effective not only for depression screening but also extending successfully to other conditions like self-harm detection, and potentially other behavioral disorders (RQ3.), across different combinations of LLMs, prompts, and retrievers (RQ2.). We also show how our approach improves upon simpler methods for screening procedure that rely on direct prompting about the presence or absence of a depressive disorder (RQ4.).
@@ 7 Limitations
Although the results are particularly promising given the available data, a limitation of this work is the relatively small number of users. Furthermore, although the method can be easily extended to other types of questionnaires, there is no guarantee that similar results will be replicated across different questionnaires or various types of mental disorders.
@@
Furthermore, in this work we use cut-off scores to define different risk thresholds for specific disorders as reported in the original works based on psychometric criteria. However, these cut-off score guidelines are typically provided with the recommendation that thresholds should be adjusted according to sample characteristics and the intended purpose of the questionnaire.
@@ 8 Ethical Considerations
Ethical considerations extend to privacy and data protection, ensuring the confidentiality and security of users social media data. In the case of this study, informed consent has been given from users for research practice, with users fully aware of how their data is being used and for what purpose.
@@ Appendix A Data Availability
Appendix B Dense Retrieval Models
We use a pool of different dense retrieval models: msmarco-MiniLM-L-6-v3333https://huggingface.co/sentence-transformers/msmarco-MiniLM-L-6-v3, msmarco-MiniLM-L-12-v3444https://huggingface.co/sentence-transformers/msmarco-MiniLM-L-12-v3, msmarco-distilbert-base-v4555https://huggingface.co/sentence-transformers/msmarco-distilbert-base-v4, sentence-T5-base [^6] https://huggingface.co/sentence-transformers/sentence-t5-base, msmarco-distilbert-base-tas-b [^7] https://huggingface.co/sentence-transformers/msmarco-distilbert-base-tas-b, all-mpnet-base-v2888https://huggingface.co/sentence-transformers/all-mpnet-base-v2, GIST [^9] https://huggingface.co/avsolatorio/GIST-Embedding-v0, sf-model-e5101010https://huggingface.co/jamesgpt1/sf_model_e5, contriever-msmarco [^11] https://huggingface.co/facebook/contriever-msmarco, bge-large [^12] https://huggingface.co/BAAI/bge-large-en. The selected models are a diverse and representative sample of the SOTA for dense retrieval and have already been tested in the literature Khramtsova et al. (2023); Barros et al. (2024); Solatorio (2024).

### v2

@@ 4.3 Proposed Framework
The proposed workflow - aRAG - involves a two-step pipeline. First, we retrieve the most relevant posts for each user with respect to each questionnaire item, and then we instruct the LLM to predict the corresponding score.
Figure 3: Results associated with the use of different combinations of LLMs, prompting strategies, and retrieval models across various metrics for the eRisk 2019 collection. The best combinations for each metric are highlighted with $\star$. The greener the color, the better the score. For readability we report only the first decimal.
Our investigation combines SOTA retrieval techniques with both open and closed-source LLMs, examining how different prompting strategies affect psychological assessment accuracy. All LLMs are used with temperature set to 0 to force deterministic outputs. For predictions, we use the closed-source models gpt-4o-mini and Claude-3.5-Sonnet, alongside large-scale open-source models Qwen 2.5 70B and DeepSeek V3, as well as lightweight open-source models Phi-3-mini and Phi-3.5-mini. For prompting, we use two techniques illustrated in Figure 2. The first technique predict the item scores directly based on relevant Reddit posts (Direct), while the second approach guides the LLM to reflect on intermediate steps (CoT), encouraging the LLM to go through reasoning steps before predicting the final score.
@@ 5 Results
eRisk 2020 ILab Martínez-Castaño et al. (2020) 27.14% 81.70% 37.07% 69.41%
@@ 5.1 Predicting Psychological Questionnaire Scores
To evaluate the effectiveness of our approach in predicting responses to the BDI-II questionnaire, we use the official eRisk benchmark metrics Losada et al. (2019) that assess performance at two distinct levels (for all the metrics considered, the higher the value, the better):
At the level of the questionnaire, we examine the Hit Rate of the Depression Category (DCHR) which measures the accuracy in estimating depression severity levels (minimal: 0-9, mild: 10-18, moderate: 19-29, and severe: 30-63), and the Average Difference between Overall Depression Levels (ADODL), which evaluates the general BDI-II score estimations.
@@
We systematically evaluate different combinations of LLMs, dense retrievers, and prompting strategies for the eRisk 2019 dataset, using the best combinations for the eRisk 2020 collection. To address (RQ.2), Figure 3 presents heatmap visualizations comparing performance across all combinations for the four evaluation metrics on the eRisk 2019 dataset.
On average, we observe that lightweight models perform significantly better with direct prompting compared to CoT approach, while this difference is less pronounced in larger architectures. Overall, we notice that closed-source LLMs and large-scale open-source models often outperform lightweight ones. This pattern can be attributed to the limited model capacity of lightweight architectures, their reduced ability to effectively manage multi-step reasoning chains during CoT prompting, and their lesser capability to capture subtle linguistic nuances crucial for mental health assessment. Specifically, the worst results are obtained when using lightweight open-source models combined with CoT prompting techniques.
After identifying the best performing combinations on the eRisk 2019 dataset from Figure 3, we use these to the 2020 collection. Table 4 shows these results alongside baseline benchmarks. For comparison, we consider the best performing models from previous work for each metric for the eRisk 2019 and 2020 collections. We refer the reader to the corresponding overview for more in-depth details Losada et al. (2019, 2020). In Section D of the Appendix we further discuss all the models used as benchmarks. Regarding the 2019 edition, the proposed approaches outperform the benchmarks across all considered metrics, except for AHR, where only one combination manages to outperform that edition’s best model. On the 2020 eRisk collection, we achieve: (1) superior item-level metrics; our approach outperforms existing benchmarks on granular metrics (AHR and ACR); (2) SOTA depression category accuracy (DCHR); we obtain the best result with a voting-regressor ensemble based on the rounded average scores of the top 3 closed-source approaches with the highest DCHR scores in 2019 evaluations (as reported in Table 4). These results are particularly notable as we maintain high performance without requiring training data, unlike all previous approaches reported that rely on the 2019 dataset for supervision. Interestingly, as shown in Table 4, our approach demonstrates consistent performance across both eRisk editions despite their different category distributions (see Table 2). Specifically, our method maintains stable performance on three key metrics (ADODL, AHR, and ACR), with only DCHR showing variation between editions. This negative result is due to how overall scores are categorized into the 4 severity categories of the BDI-II questionnaire. Within the context of the challenge, the authors of the eRisk workshop use the BDI parameterization, the version preceding BDI-II. In BDI-II, new ranges are introduced that change from those of the previous test, especially regarding the minimum level of depression. Specifically, minimal or absent depression is identified as 0-13, Mild as 14-19, Moderate as 20-28, and severe as 29-63 Beck (1996); Warmenhoven et al. (2012). Table 5 shows how DCHR changes when using the correct reparameterization, obtaining excellent and completely counterintuitive results, especially for the two models using Claude and the CoT strategy, compared to those obtained with the previous questionnaire parameterization in the 2020 dataset. In Section E.1 of the Appendix, we further justify our aRAG approach’s effectiveness by comparing it against non-RAG baselines, where we test closed-source LLMs (gpt-4o-mini and Claude-3.5-Sonnet) using direct input of all posts within their context window.
@@
Figure 4: Performance of our approach compared to the best approach in each of the two eRisk editions for the detection of self-harm, Anorexia, and Pathological Gambling. Models with use training data while models with are fully unsupervised. Precision, Recall, and F1 metrics are reported.
5.2 Beyond Depression: Identifying Signs of different MHCs through questionnaire
In this section, as a proof-of-concept, we extend the use of questionnaires to different mental health conditions, such as self-harm, anorexia, and pathological gambling. We approach the identification of self-harm behaviors using the Self-Harm Inventory (SHI), a 22-item, yes/no self-report questionnaire, that screens for lifetime history of self-harm behaviors. A score of 5 or more “yes” responses on the SHI indicates potential mild forms of Deliberate Self-Harm (DSH) Latimer et al. (2009). To answer (RQ.3), the methodology follows the same approach used for BDI-II, with the key distinction that SHI is structured as a binary questionnaire, unlike BDI-II Likert scale. The process involves retrieving the most relevant posts using SHI questions as queries, followed by LLM-generated responses. For SHI, since each item has binary yes/no responses, we use only the questions as retrieval queries. For this specific task, we chose the combination of gpt-4o-mini with Direct Prompt and MiniLM-L12-v3 as the retrieval model, as this configuration demonstrated the best trade-off between performance, computational costs, and processing speed in previous tests. The dataset provides each user complete post history, associated with a binary label indicating the presence or absence of self-harm behaviors. While the original task aims to identify self-harm cases as early as possible using the minimum number of posts, our approach considers the entire history to maximize prediction accuracy. For the 2019 edition, we compared our approach with two reference models: UNSL Burdisso et al. (2019), which analyzes only a subset of posts, and iLab Martínez-Castaño et al. (2020), which uses BERT fine-tuned in a fully supervised context. iLab approach was optimized to maximize F1-score and trained on a custom dataset built from self-harm subreddit posts. To ensure a fair comparison, we considered the version of iLab that has access to the complete post history for both the 2019 and 2020 editions. Figure 4 shows that while the fine-tuned BERT-based model, benefiting from an extensive training corpus, generally outperforms our unsupervised approach, the performance gap is remarkably narrow, particularly for the 2019 edition. Notably, our approach achieves superior precision in the 2020 edition. These results are particularly significant considering that our approach do not need any training data and requires no training data, yet achieves competitive performance compared to fully supervised models that leverage extensive domain-specific training.
We further applied the same methodology to the early detection of anorexia and pathological gambling, using the SCOFF questionnaire and the Pathological Gambling Diagnostic Form respectively. In both cases, we used the same pipeline as in the SHI scenario. As reported in Figure 4, our method achieves the highest F1 score for anorexia and gambling tasks, outperforming the top-scoring systems of the respective editions Mohammadi et al. (2019); Ragheb et al. (2019); Mármol-Romero et al. (2022); Fabregat et al. (2022) in this specific metric. These findings reinforce the generalizability of our adaptive RAG framework, demonstrating that standardized psychological questionnaires can effectively guide LLMs to detect a wide range of MHCs in a fully unsupervised setting. This suggests that our unsupervised approach based on adaptive RAG can offer a viable alternative in scenarios where labeled training data is scarce or unavailable.
Even though in some cases the results do not outperform the benchmark models, this may be attributed not only to the lack of training data in our approach, but also to the fact that our questionnaire-guided approach occasionally fails to find sufficient evidence to support a diagnosis based on the psychometric tool used. As a result, our approach tends to be more conservative in assigning final diagnoses, sometimes refraining from matching the diagnostic criteria defined by the questionnaires unless the retrieved content provides clear and consistent indicators.
5.3 The Importance of Questionnaire for Screening Procedure
Model $\mathbf{\tau}$ P R F1
@@ 5.3 The Importance of Questionnaire for Screening Procedure
For this task, we used data from the 2017 eRisk edition to address (RQ4.). As in the previous subsection, we have access to users post histories along with corresponding labels indicating whether each user exhibited clinical signs of depression. To prevent trivial classification by the LLM, we removed the word depression and related terms from posts in which users explicitly self-diagnosed.
@@
We tested the same aRAG approach that proved effective for previous tasks, using the same combination of methods but guided by different types of standardized depression screening questionnaires. Specifically, we compared results obtained from the Beck Depression Inventory-II (BDI-II), Patient Health Questionnaire-9 (PHQ-9), Depression Anxiety Stress Scales-42 (DASS-42) (specifically focusing on its depression subscale), and simply instructing the LLM to determine if the patient was suffering from depression. While for BDI-II we used choice options as queries, for PHQ-9 and DASS-42 we only used the questionnaire items as queries since only the questions contain textual content.
@@
Each of the three questionnaires includes an established optimal cut-off score, denoted as $\tau$ in Table 6, for identifying clinically significant depression, specifically at the moderate severity threshold. The table also reports the results obtained from the different approaches. The worst results were obtained through direct instruction to gpt-4o-mini and through the use of the DASS-42 depression subscale. The best results in terms of precision were achieved using PHQ-9 (0.69), while BDI-II showed the highest recall (0.76). The highest F1 score (0.60) was achieved using an ensemble classifier combining all three questionnaires. To answer (RQ4.), our findings suggest that structured psychological assessment tools, can enhance the effectiveness of LLM-based mental health assessments compared to direct questioning approaches.
@@ 6 Conclusions
We introduce a novel aRAG approach that leverages standardized psychological questionnaires to guide LLMs in mental health screening, requiring no training data. We demonstrate the advantage of our approach in supporting mental health screening tasks. The results show the advantages of our approach in: automatically completing questionnaires (RQ1.), proving effective not only for depression screening but also extending successfully to other conditions like self-harm detection, anorexia, and pathological gambling, as well as potentially other MHCs (RQ3.), across different combinations of LLMs, prompts, and retrievers (RQ2.). We also show how our approach improves upon simpler methods for screening procedure that rely on direct prompting about the presence or absence of a depressive disorder (RQ4.), by providing a structured interpretation of the user’s psychological state and enabling an estimation of the severity level through clinically questionnaire scores.
@@ 7 Limitations
Although the results are particularly promising given the available data, a limitation of this work is the relatively small number of users. Furthermore, although the method can be easily extended to other types of questionnaires, there is no guarantee that similar results will be replicated across different questionnaires or various types of MHCs.
@@
Furthermore, in this work we use cut-off scores to define different risk thresholds for specific disorders as reported in the original works based on psychometric criteria. However, these cut-off score guidelines are typically provided with the recommendation that thresholds should be adjusted according to sample characteristics and the intended purpose of the questionnaire. Additionally, such cut-offs may not be fully consistent in the context of social media analysis and may require further adaptation.
@@ 8 Ethical Considerations
Ethical considerations extend to privacy and data protection, ensuring the confidentiality and security of users social media data.
@@ Appendix A Data Availability
Appendix B ABIDE
Adaptive Binomial Intrinsic Dimension Estimator (ABIDE) Noia et al. (2024) is an approach to estimate the Intrinsic Dimension (ID) of the data. The idea of ID is to quantify the complexity of high-dimensional datasets. In essence, it represents the minimum number of variables needed to describe the underlying structure of data without significant loss of information Denti et al. (2022, 2023). In fact, in many real-world scenarios, especially for what concerns text, there are often hidden relationships and dependencies among these features. This means that the data might actually lie on a lower-dimensional manifold embedded within the high-dimensional space.
However, estimating ID can be difficult, especially in real-world scenarios, where datasets often have complex structures and ID can vary depending on the scale at which the data are observed.
ABIDE addresses the scale-dependency challenge through a novel adaptive approach. At its core, ABIDE uses the concept of $k^{*}$, which represents the optimal neighborhood size for each data point. $k^{*}$ is not fixed across all the observations, but varies for each point, allowing the method to adapt to local data characteristics. The algorithm works iteratively. It starts with an initial ID estimate, using the 2NN method Facco et al. (2017). Then, for each data point, it determines the largest neighborhood ($k^{*}$) where the data density can be considered approximately constant. Using these $k^{*}$ values, ABIDE recalculates the ID estimate. This process is repeated, refining at each iteration both the ID estimate and the $k^{*}$ values for each point. The iteration continues until convergence is reached.
Appendix C Dense Retrieval Models
We use a pool of different dense retrieval models: msmarco-MiniLM-L-6-v3222https://huggingface.co/sentence-transformers/msmarco-MiniLM-L-6-v3, msmarco-MiniLM-L-12-v3333https://huggingface.co/sentence-transformers/msmarco-MiniLM-L-12-v3, msmarco-distilbert-base-v4444https://huggingface.co/sentence-transformers/msmarco-distilbert-base-v4, sentence-T5-base [^5] https://huggingface.co/sentence-transformers/sentence-t5-base, msmarco-distilbert-base-tas-b [^6] https://huggingface.co/sentence-transformers/msmarco-distilbert-base-tas-b, all-mpnet-base-v2777https://huggingface.co/sentence-transformers/all-mpnet-base-v2, GIST [^8] https://huggingface.co/avsolatorio/GIST-Embedding-v0, sf-model-e5999https://huggingface.co/jamesgpt1/sf_model_e5, contriever-msmarco [^10] https://huggingface.co/facebook/contriever-msmarco, bge-large [^11] https://huggingface.co/BAAI/bge-large-en. The selected models are a diverse and representative sample of the SOTA for dense retrieval and have already been tested in the literature Khramtsova et al. (2023); Barros et al. (2024); Solatorio (2024).